# Restless Bandit Problem with Rewards Generated by a Linear Gaussian Dynamical System

### v1

@@ C.1 Model Error
As it is shown in Lemma 2, inequality (22) is satisfied with a probability of at least $1-\delta$. For the term $\varepsilon_{a;\tau}$, since $\varepsilon_{a;\tau}$ is conditionally $B_{R}$-sub-Gaussian on $\mathcal{F}_{t-1}$ and $\Xi_{t}$ is $\mathcal{F}_{t-1}$ measurable, then the conditions described in [17], Theorem 1, are satisfied. Therefore, the following inequality is satisfied with a probability of at least $1-\delta$:
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}.$
@@
Theorem 3 states that with a probability of at least $\left(1-\delta\right)^{2}$, the weighted-norm in (29) is upper-bounded by two terms, (14) and (15). Note that the terms $\delta$ are within the interval $(0,1)$, where the closer $\delta$ is to 00, the larger the terms (14) and (15) are. The opposite occurs if $\delta$ are closer to $1$.
@@ C.2 Prediction Error
With a probability of at least $\left(1-\delta\right)^{2}$, the following inequality is satisfied
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\leq\sigma\left(\delta,c_{a}\mid\mathbf{c}\right),$ (31)
where $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is defined to be
$\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)\triangleq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (32)
@@
$\begin{array}[]{cc}\underset{a\in[k]}{\arg\max}&\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ \mbox{s.t.}&\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\end{array}.$ (33)
@@
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right),$
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ -\frac{\alpha}{2}\left(\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right),$ (34)
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\\ +\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}-\frac{\alpha}{2}\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}+\frac{\alpha}{2}\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}.$ (35)
@@
$\frac{\partial L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)}{\partial\alpha}$ $=-\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right).$
@@
$\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\left\|\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0,$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}$ $=\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}$ $=\alpha^{2},$
$\Rightarrow\alpha=\sqrt{\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}}.$ (37)
@@
$\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)=\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)\sqrt{\frac{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}}.$
@@
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}.$ (38)
If we define $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ to be (32), then the inequality (31) is satisfied with a probability of at least $\left(1-\delta\right)^{2}$. ∎
@@
The upper bound of $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is satisfied with a probability of at least $1-\delta$ where the bound is
$\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)\leq B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right),$ (39)
where $B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)$ is defined to be
$B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(s\lambda+n\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\\ +\sqrt{n}B_{c}B_{R}\sqrt{s}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}+\lambda B_{G}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (40)
@@
The upper bound of $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is as follows: given that the following inequality is satisfied with a probability of least $1-\delta$:
@@
then according to Lemma 11 in [17], $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ is upper bounded as (with a probability of at least $1-\delta$)
$e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(\mbox{tr}\left(\lambda I_{s}\right)+\left|\mathcal{T}_{c_{a}\mid\mathbf{c}}\right|\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\sqrt{\det\left(\lambda I_{s}\right)}}\right)}$
$\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(s\lambda+n\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}.$
The term $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ is based on bound (22). To bound $\sqrt{\mbox{tr}\left(I_{s}-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}$ in (22), note that since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$, then
@@
Therefore, we have the upper bound of $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$:
$b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\leq\sqrt{n}B_{c}B_{R}\sqrt{s}.$
Finally, the term $\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}$ has the following bound with a probability of at least $1-\delta$
$\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}$ $\leq\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}$
$\leq\sqrt{\mbox{tr}\left(\frac{I_{s}}{\lambda}\right)}\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2},$
$\Rightarrow\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}\leq\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (42)
where the final inequality (42) uses (41). Therefore, the bound for $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is (39), which satisfied with a probability of at least $1-\delta$. ∎

### v2

@@ C.1 Model Error
As it is shown in Lemma 2, inequality (22) is satisfied with a probability of at least $1-\delta$. For the term $\varepsilon_{a;\tau}$, since $\varepsilon_{a;\tau}$ is conditionally $B_{R}$-sub-Gaussian on $\mathcal{F}_{t-1}$ and $\Xi_{t}$ is $\mathcal{F}_{t-1}$ measurable, then the conditions described in [17], Theorem 1, are satisfied. Therefore, the following inequality is satisfied with a probability of at least $1-\delta_{e}$:
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta_{e}}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}.$
@@
Theorem 3 states that with a probability of at least $\left(1-\delta_{b}\right)\left(1-\delta_{e}\right)$, the weighted-norm in (29) is upper-bounded by two terms, (14) and (15). Note that the terms $(\delta_{e},\delta_{b})$ are within the interval $(0,1)$, where the closer $(\delta_{e},\delta_{b})$ are to 00, the larger the terms (14) and (15) are. The opposite occurs if $(\delta_{e},\delta_{b})$ are closer to $1$.
@@ C.2 Prediction Error
With a probability of at least $\left(1-\delta_{e}\right)\left(1-\delta_{b}\right)$, the following inequality is satisfied
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\leq\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right),$ (31)
where $\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)$ is defined to be
$\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)\triangleq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (32)
@@
$\begin{array}[]{cc}\underset{a\in[k]}{\arg\max}&\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ \mbox{s.t.}&\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\end{array}.$ (33)
@@
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right),$
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ -\frac{\alpha}{2}\left(\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right),$ (34)
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\\ +\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}-\frac{\alpha}{2}\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}+\frac{\alpha}{2}\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}.$ (35)
@@
$\frac{\partial L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)}{\partial\alpha}$ $=-\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right).$
@@
$\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\left\|\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}\right)$ $=0,$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}$ $=\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}}$ $=\alpha^{2},$
$\Rightarrow\alpha=\sqrt{\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}}}.$ (37)
@@
$\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)=\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)\sqrt{\frac{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)^{2}}{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}}.$
@@
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}.$ (38)
If we define $\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)$ to be (32), then the inequality (31) is satisfied with a probability of at least $\left(1-\delta_{e}\right)\left(1-\delta_{b}\right)$. ∎
@@
The upper bound of $\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)$ is satisfied with a probability of at least $1-\delta$ where the bound is
$\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)\leq B\left(\delta\mid\mathbf{c}\right),$ (39)
where $B\left(\delta\mid\mathbf{c}\right)$ is defined to be (17).
@@
The upper bound of $\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)$ is as follows: given that the following inequality is satisfied with a probability of least $1-\delta$:
@@
then according to Lemma 11 in [17], $e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)$ is upper bounded as (with a probability of at least $1-\delta$)
$e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)$ $\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta_{e}}\frac{\left(\mbox{tr}\left(\lambda I_{s}\right)+\left|\mathcal{T}_{c_{a}\mid\mathbf{c}}\right|\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\sqrt{\det\left(\lambda I_{s}\right)}}\right)}$
$\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta_{e}}\frac{\left(s\lambda+(n-s)\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}.$
The term $b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)$ is based on bound (22). To bound $\sqrt{\mbox{tr}\left(I_{s}-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}$ in (22), note that since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$, then
@@
Therefore, we have the upper bound of $b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)$:
$b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\leq\sqrt{n-s}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{s}.$
Finally, the term $\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$ has the following bound with a probability of at least $1-\delta$
$\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$ $\leq\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}$
$\leq\sqrt{\mbox{tr}\left(\frac{I_{s}}{\lambda}\right)}\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2},$
$\Rightarrow\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}\leq\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (41)
where the final inequality (41) uses (40). Therefore, the bound for $\sigma\left(\delta_{e},\delta_{b},c_{a}\mid\mathbf{c}\right)$ is (39), which satisfied with a probability of at least $1-\delta$. ∎